3^1-2y=square root of 3
Answer by stanbon(75887)

3^(1-2y) = sqrt(3)
----
3^(1-2y) = 3^(1/2)
1-2y = 1/2
2y = 1-(1/2)
2y = 1/2
y = 1/4
============
Cheers
Stan H.